### v3

# ChOiRe: Characterizing and Predicting Human Opinions with Chain of Opinion Reasoning

Abstract
Warning: This paper includes examples that may be deemed sensitive or offensive.
Aligning language models (LMs) with human opinion is challenging yet vital to enhance their grasp of human values, preferences, and beliefs. We present ChOiRe [^1] Codes and data will be available at https://github.com/dxlong2000/ChOiRe., a four-step framework to predict human opinion which differentially models the user’s explicit personae (i.e. demographic or ideological attributes) that are manually declared, and implicit personae inferred from user historical opinions. ChOiRe consists of (i) an LM analyzing the user’s explicit personae to filter out irrelevant attributes; (ii) the LM ranking the implicit persona opinions into a preferential list; (iii) Chain-of-Opinion (CoO) reasoning, where the LM sequentially analyzes the explicit personae and the most relevant implicit personae to perform opinion prediction; (iv) and where ChOiRe executes Step (iii)’s CoO multiple times with increasingly larger lists of implicit personae to overcome insufficient personae information to infer a final result. ChOiRe achieves new state-of-the-art effectiveness with limited inference calls, improving previous techniques significantly by 3.22%. We also show that ChOiRe’s Steps (i) and (ii) can significantly better fine-tune opinion-aligned models, by up to 18.44%.
1 Introduction
Language models (LMs) are becoming indispensable tools, serving in various roles such as dialogue agents OpenAI (2022); Google (2022), data analysts Wang et al. (2023a); Cheng et al. (2023), and decision support Ye et al. (2023). LMs also demonstrate the capability to model distinct opinions which influence response generation on input queries Bai et al. (2022); Glaese et al. (2022); Santurkar et al. (2023). Unfortunately, the opinions modeled by language models are shaped by the extensive training and feedback data, which are themselves influenced by countless human perspectives, making them inherently challenging to model. As human–AI interactions become common, it becomes imperative to align models with human opinion to meet individual and personalized expectations.
Alignment frameworks such as RLHF Christiano et al. (2017); Ouyang et al. (2022) have been devised to train and fine-tune LMs. Nonetheless, applying them to align large language models (LLMs) with human opinions is difficult because they require both significant computational resources and high-quality supervised feedback data, which is itself challenging to collect. Therefore, prompt-based opinion alignment through persona has been studied as an easy-to-deploy and easy-to-combine resource-parsimonious alternative Perez et al. (2023); Simmons (2023); Santurkar et al. (2023); Deshpande et al. (2023).
However, even when aligning with well-represented groups, persona-based methods result in low steerability Santurkar et al. (2023). This raises significant concerns when using LMs to model for individual users. Since individuals hold nuanced opinions that evolve and are influenced by situational factors, it is more challenging to achieve individual opinion LM alignment. Hwang et al. (2023) find significant opinion variations among individuals sharing the same demographics, exposing flaws in current group-focused alignment. They argue for individualized models, introducing an approach (detailed below) integrating user’s demographic and ideological attributes, (we term as explicit personae) and historical opinions (implicit personae) into the prompt for opinion prediction.
While Hwang et al. (2023)’s naïve integration achieves promising results, we argue that it suffers from a few key limitations. First, it employs all explicit personae. However, we contend that only a subset is needed for accurate opinion prediction; including non-relevant personae may act as noise, harming predictive performance. Second, Hwang et al. (2023) utilize the top-$K$ semantically similar opinions with the question (here termed top-$K$ implicit personae). This approach is inefficient, as such similar opinions may not offer the most valuable information for opinion prediction. Additionally, our empirical experiments suggest that LMs may lack sufficient personae evidence with a fixed $K$ — dynamically adjusting $K$ per task can overcome such deficiencies. Finally, while Chain-of-Thought (CoT; Wei et al. 2022; Kojima et al. 2022) enables LMs to perform multi-step reasoning tasks effectively, we find that the naïve application of CoT does not help modern LLMs like ChatGPT with opinion alignment [^2] We validate all our claims on these limitations in Sections E.1, 6.1, 4 and 5, respectively..
To address the above challenges, we propose ChOiRe [^3] Chain of Opinion Reasoning, pronounced as the English word “choir”.(fig. 1), a four-step solution for opinion prediction leveraging LLMs’ strong data evaluation and analytic capabilities Wang et al. (2023a); Cheng et al. (2023). First, an LLM analyzes a target user’s explicit personae to discard irrelevant ones. Second, the LLM ranks implicit persona opinions in order of usefulness, selecting the top-$K$ as the most valuable. This surpasses the constraint of using semantic similarity scores. Third, we introduce Chain-of-Opinion (CoO), a designed variant of CoT that allows the LLM to explain and analyze selected explicit personae and top-$K$ implicit personae sequentially. ChOiRe applies self-consistency over CoO to provision the appropriate amount of user information for opinion inference.
ChOiRe achieves new state-of-the-art (SOTA) in opinion alignment effectiveness and reliability, while using a limited inference budget. We conduct a thorough analysis to verify our hypotheses concerning explicit and implicit personae and defend our Chain-of-Opinion reasoning methodology. Our analysis further reveals that ChOiRe’s first two steps in handling explicit and implicit personae also helps to fine-tune opinion-aligned models.
2 Related Work
Aligning LMs with Humans.
Aligning language models with human behaviour is a recent area of study as alignment can increase user experience satisfaction and utility Wang et al. (2023c). One line of work develops prompting techniques with user demographic information (e.g., political identity) to encourage LMs to output human-like responses. Argyle et al. (2023) show that by properly conditioning LMs with targeted identity and personality profiles, it is possible to produce biased outputs that strongly correlate with human responses. Furthermore, Simmons (2023) claims that LLMs are moral mimics: by giving models a political identity, they produce texts mirroring the associated moral biases. Despite recent advances, Santurkar et al. (2023) discovered that LMs align poorly with human opinions, as evidenced by model performance on public opinion polls. Hwang et al. (2023) recently propose to incorporate explicit and implicit personae to predict human opinions in new contexts. In section 1, we argue that this naïve strategy is suboptimal. ChOiRe overcomes these limitations.
Reasoning with LMs via Prompting.
Large-scale model architectures Devlin et al. (2019); Radford et al. (2019); Brown et al. (2020); Chowdhery et al. (2023); Touvron et al. (2023) have enabled large language models (LLMs) to excel at various NLP tasks using zero- or few-shot prompting Liu et al. (2023). Notably, Wei et al. (2022); Kojima et al. (2022) propose prominent Chain-of-Thought (CoT) techniques, enabling LLMs to explicate intermediate reasoning steps to solve multi-step reasoning tasks with higher fidelity and efficiency.
Can CoT analyze and predict human opinion effectively? We find that a naïve application of CoT does not help GPT-X models (section 5), but that an appropriate modification does. We propose Chain-of-Opinion (CoO) reasoning (section 3) that overcomes CoT’s limitations in this task. Noting that prompting techniques such as task decomposition Khot et al. (2023); Zhou et al. (2023) and retrieved-based methods Yao et al. (2023); Shinn et al. (2023) have been recently introduced, we focus only on the reasoning explanation aspect here given the abstractive and challenging nature of the task.
Figure 1: ChOiRe overview, consisting of the four main steps (cyan background), as detailed in section 3.
3 ChOiRe: A Chain of Opinion Framework
Task Formalisation. We follow Santurkar et al. (2023), and formulate the opinion prediction task as multiple-choice question answering. Formally, a benchmark with $N$ data points is denotated as $D=\{\langle T,E,I,q,a\rangle_{n}\}_{n=1}^{N}$, where $T$, $E$ and $I$ indicate the (T)opic of a question $q$, the (E)xplicit personae and (I)mplicit personae of the user answering $q$ with opinion $a$. Following the prior work, $E$ consists of $12$ user demographic and ideology metadata attributes, and $I$ contains a number of the user’s historical opinions in the format of question–answer pairs. Models then learn to analyze $T,E,I,q$ and predict the opinion $a$.
Fig. 1 shows an overview of ChOiRe, consisting of four main steps (marked with a cyan background). First, ChOiRe employs an LLM to analyze and select a subset of relevant explicit personae, denoted as $E^{rel}\subseteq E$ for answering the opinion question $q$. The LLM then assesses the informativeness of the implicit personae ($I$) in predicting $q$, selecting the top-$K$ implicit personae (termed LLMtop-$K$). Next, an LLM is prompted to explain the provided explicit $E^{rel}$ and implicit LLMtop-$K$ personae sequentially in a Chain-of-Opinion (CoO) reasoning strategy. Finally, ChOiRe calls the LLM to predict the opinion $a$ with varying values of $K$ for the top-$K$ implicit personae. ChOiRe chooses the opinion with the highest frequency as the final prediction We detail the steps below.
3.1 Filtering Explicit Personae Attributes (FEA)
Accounting for explicit personae, which consist of the demographic and ideological metadata attributes of users — such as their age, income, and political ideology — has been shown to help models characterize and predict human opinions more accurately Hwang et al. (2023). However, which personae matter and which do not? are still open questions. Section E.1 shows such an example in full, where, when considering all of explicit personae, the model makes an incorrect prediction while removing unnecessary personae the model made a correct prediction. This may be caused by the LLM’s attention mechanism’s forcing the model to attend to all input tokens, even irrelevant ones. To filter out unnecessary explicit personae, we ask the LLM to reason and analyze how each persona is helpful for the model to predict the opinion via Chain-of-Thought to output a list of which personae are relevant given the question and the opinion answer choices [^4] We provide our FEA prompt in Section C.1. Surprisingly, we find that LLMs evaluate more than half of the explicit personae as irrelevant on average. We further conduct human evaluations to verify this finding in section 4.
3.2 Implicit Personae Opinions Ranking (LLMtop-$K$)
LLMs have been sensitive to selected demonstrations and their order in the prompts Perez et al. (2021); Luo et al. (2023); Gao et al. (2023). For predicting human opinions, we discover that LLMs are also sensitive to the chosen implicit personae opinions as input. Hwang et al. (2023) rank the implicit personae opinions via semantic-similarity scores and select top-$K$. This strategy is suboptimal because the top-ranked opinions in terms of semantic similarity may not be the ones that provide the most supportive information for the models to predict opinions (section E.2). As LLMs are shown to be good data analysts Wang et al. (2023a); Cheng et al. (2023), we propose to address the above challenge by utilizing LLMs to analyze and rank the implicit personae opinions in usefulness descending order. Our finding is that despite the output rankings from LLMs varying with different input orders of implicit personae opinions, the sets of LLMtop-$K$ opinions overlap by a good coefficient when $K$ is large enough ($\geq 8$) (fig. 5). Therefore, we propose to input the implicit personae opinions to LLMs in a random order to make our method more versatile. We also examine the case where we input the opinions in the semantic similarity order. We illustrate the prompt template in section C.2. By performing this step, our proposed method supports the usefulness of opinions in predicting the test opinions, rather than the semantic similarity. We term this method as LLMtop-$K$.
3.3 Chain-of-Opinion Reasoning (CoO)
Wei et al. (2022); Kojima et al. (2022) introduce few-shot and zero-shot Chain-of-Thought (CoT) prompting strategies demonstrating that by reasoning step-by-step, LLMs can achieve promising results on complex tasks. However, the sampled reasoning steps can be inconsistent, leading to possibly different outcomes Wang et al. (2023b). Furthermore, it is little known how the models perceive multiple implicit personae opinions, especially when many opinions are input, which one(s) the models used, which one(s) they didn’t for predicting the opinion? Our preliminary experiments with CoT (sections 6.1, E.4 and E.3) reveal that the CoT explanations can vary frequently based on different subsets of opinions mentioned in their explanations, leading to diverse final answers, especially when the decoding temperature is relatively high555greater or equal to 0.6. To mitigate this issue, we propose to instruct the LLMs to analyze the given explicit and implicit personae one by one before concluding the prediction via simply adding "explaining and analyzing how each of the Opinions and Demographic Information supports the question" into the prompt instruction. Given an LLM that can follow human instructions well such as ChatGPT OpenAI (2022), this addition offers two notable advantages. First, for each question, we ensure that the model explains and analyzes the provided personae one by one without missing any, possibly resulting in more thorough predictions. Second, this method helps the model to output more consistent reasoning explanations, enhancing its reliability (section 6.1).
3.4 Answer Consistency with Dynamic Numbers of Opinions
Prior work Hwang et al. (2023) fixes the number of implicit personae opinions for prediction to $K=8$. However, this approach occasionally results in models generating "...the answer cannot be determined." (tables 4 and E.5). We attribute this to insufficient user implicit personae opinions provided. Inspired by Self-Consistency (SC) Wang et al. (2023b), our approach involves sampling multiple answers using different $K$ values for a given question. The most frequent answer, along with its explanation, becomes the final prediction. Our method is distinct from SC since SC samples multiple answers with a fixed prompt. We experiment with $K\in\{8,10,12\}$ for efficiency.
4 Evaluation
Dataset.
We experiment on OpinionQA dataset Santurkar et al. (2023) — the only opinion QA dataset to date consisting of both user explicit and implicit personae designed for the assessment of alignment between LMMs’ opinions and human participants, encompassing a diverse range of $60$ US demographic groups. It covers $60$ US demographic groups, with $15$ topics, each comprising around $100$ questions, gathered from $5,340$ users.
Dataset Preprocessing.
Due to limited resources, we randomly sample $25$ users per topic for our experiments. For each user, we follow Hwang et al. (2023) to use $20\%$ of the implicit questions as the implicit persona. For the remaining $80\%$ implicit questions, we randomly select a maximum of $15$ implicit questions for testing. Our sampling method results in a total of $375$ users and $5,603$ implicit evaluation question–answer pairs. Our subset is highly representative because we gather 25 users from every topic and 15 questions per user. Rigorous statistical tests further validate the significance of our results which align closely with Hwang et al. (2023) testing on a larger subset using InstructGPT.
Prompting Baselines.
We use both closed-source ChatGPT OpenAI (2022), ChatGPT-Instruct OpenAI (2023a), GPT-4 OpenAI (2023b), and open-source Mistral-7B-Instruct-v.02 Jiang et al. (2023) as our LLMs, and compare ChOiRe with $5$ prompting methods: (1) W/o persona, where LLMs are evaluated without user historical opinions, ideology, or demographic data; (2) Demographic + Ideology + top8 Opinions (termed DIO-top8), introduced by Hwang et al. (2023) demonstrating that integrating explicit and implicit personae enhances user opinion modeling and prediction, achieving state-of-the-art results on OpinionQA at that time; (3) DIO-top8 + CoT is the Chain-of-Thought (CoT) prompting Kojima et al. (2022) version of DIO-top8 involving appending "answer the following question step-by-step" to prompts, aiming to explore whether CoT improves model performance in this task; (4) DIO-top8 + SC is the baseline which we apply the Self-Consistency technique with CoT Wang et al. (2023b) to DIO-top8 to select the most frequent answer generated by the model as the final opinion prediction; (5) DIO-top8 + Self-refine Madaan et al. (2023) interactively feedbacks and refines the answers by LLMs. We do not experiment with InstructGPT Ouyang et al. (2022) like Hwang et al. (2023) since this model is going to be deprecated and replaced by ChatGPT-Instruct. For GPT-4, we only run the main experiment and we use ChatGPT for FEA and LLMtop-$K$ steps due to our limited budget. All the prompts and costs are in appendix C, implementation details in section A.1, and more baselines in section B.1.
Fine-tuning Baselines.
We further investigate whether ChOiRe’s FEA and LLMtop-$K$ steps (section 3) also improve fine-tuning for opinion-aligned models. We first create the fine-tuning data by using ChatGPT to perform ChOiRe’s FEA and LLMtop-$K$ steps ($K=8$) on a training set of $30,000$ samples randomly selected from OpinionQA which are different from our $5,603$ test ones. We then fine-tune and evaluate GPT-2 models (base, large) Radford et al. (2019) and FlanT5 models (base, large) Chung et al. (2022). Fine-tuning details are provided in section A.1.
Metrics.
We employ Accuracy and Collapsed Accuracy as the automatic evaluation metrics following Hwang et al. (2023). It is worth noting that Precision/Recall/F1 is not applicable in our task, since the numbers of answer choices are not the same for all the OpinionQA samples. In addition, human evaluations are crucial due to the absence of automated metrics assessing LLMs’ performance in FEA, LLMtop-$K$ and CoO steps of ChOiRe. Therefore, we conduct our human assessments to address these research questions: (1) LLMs’ effectiveness in filtering unnecessary explicit personae; (2) LLMs’ proficiency in ranking implicit personae opinions; (3) LLMs’ ability to explain answers via CoO. To this end, we randomly select $100$ answers generated by ChOiRe with ChatGPT, ChatGPT-Instruct, GPT-4, Mistral. We then hire $3$ excellent undergraduates who are native English speakers as annotators. For FEA and LLMtop-$K$ steps, each annotator is instructed to rate on a 1-3 scale (3 is best) via the Satisfaction criterion defined as how well the algorithm of LLMs performs in filtering/ranking, subjectively. To answer (3), we use two criteria named Reasonableness measuring how well the LLMs reason with the CoO explanations, and Follow the Instruction assessing the capability of LLMs in following our instruction to explain and predict the opinions. Three annotators are also guided to rate the criteria on a 1-3 scale. Each metric’s final score is the average of three annotators’ scores. The scoring instructions are in section D.1 and the inter-annotators’ agreement is assessed by Kripp’s alpha Krippendorff (2011).
5 Main Results
Model ChatGPT ChatGPT-Inst GPT-4 Mistral-7B-Ins.-v0.2
W/o persona 46.60/65.72 44.91/63.60 - 41.24/59.54
DIO-top8 50.22/69.21 51.95/71.16 57.98/76.86 44.16/62.47
DIO-top8 + Self-refine 43.14/65.33 42.71/62.98 - 36.23/55.06
DIO-top8 + CoT 49.96/69.05 51.90/71.51 - 52.25/71.95
DIO-top8 + SC 50.58/69.66 52.06/71.87 - 53.14/72.88
DIO-top8 + FEA 50.64/69.85 52.63/72.30 - 44.99/64.09
DIO-top8 + CoO 50.97/70.22 52.08/71.65 - 53.79/73.59
DIO-LLMtop8 51.03/70.31 52.80/72.60 - 45.86/64.98
DIO-LLMtop8 + FEA 51.19/70.69 52.97/72.84 - 45.23/64.73
DIO-LLMtop8 + FEA + CoO 51.90/71.57 53.01/72.91 59.02/78.70 54.21/74.09
ChOiRe 52.21normal-$\dagger$/72.09normal-$\dagger$ 53.26normal-$\dagger$/73.26normal-$\dagger$ 59.30normal-$\dagger$/78.82normal-$\dagger$ 54.43normal-$\dagger$/74.34normal-$\dagger$
% Improvements +3.22/+3.49 +2.52/+1.93 +2.28/+2.55 +2.42/+2.00
Table 1: Overall Accuracy/Collapsed Accuracy on ChatGPT, ChatGPT-Instruct, and GPT-4. FEA is our first step, Filtering Explicit Attributes. LLMtop8 is the second step, ranking and selecting top-8 implicit persona opinions as demonstrations, CoO is Chain-of-Opinion reasoning. Improvements are calculated with the best baseline. $\dagger$ denotes our model outperforms baselines significantly with p-value < 0.01 under t-test.
Model GPT-2-base GPT-2-large FlanT5-base FlanT5-large
W/o persona 41.14/58.87 21.94/39.11 48.98/68.33 39.83/58.43
DIO-top8 21.23/38.64 24.94/42.22 55.00/74.98 54.94/74.79
DIO-top8 + FEA 22.62/40.97 25.65/45.21 55.78/75.34 58.77/77.26
DIO-LLMtop8 22.65/41.12 28.86/47.60 57.97normal-$\dagger$/77.46normal-$\dagger$ 58.20/77.56
DIO-LLMtop8 + FEA 25.05/44.41 29.54normal-$\dagger$/48.66normal-$\dagger$ 57.45/77.13 59.00normal-$\dagger$/78.46normal-$\dagger$
% Imp. over DIO-top8 +17.99/+14.93 +18.44/+15.25 +5.40/+3.30 +7.38/+4.90
Table 2: Performance of fine-tuned baselines with our proposed FEA and LLMtop8 steps preprocessed by ChatGPT. $\dagger$ denotes our model significantly outperforms baselines with p-value < 0.01 under t-test.
Overall Prompting Results.
Table 1 shows our macro experimental outcomes. We derive $4$ main observations in this task. First, ChOiRe improves the best among baselines significantly with $3.22\%$, $2.52\%$, $2.28\%$, $2.42\%$ accuracy for ChatGPT, ChatGPT-Instruct, GPT-4, and Mistral. It establishes a strong SOTA result with GPT-4, surpassing previous SOTA DIO-top8 with InstructGPT achieving 53.74% Hwang et al. (2023) by a notable margin. Notably, in the case of GPT-4, we utilize ChatGPT for FEA and LLMtop-$K$ steps, showcasing the strength of a weaker model that enhances a stronger one. Second, we see that Accuracy and Collapsed Accuracy have the same trend, and ChOiRe also achieves the SOTA on Collapsed Accuracy with the highest improvement of $3.94\%$ observed with ChatGPT. Third, naïve CoT ("answer the following question step-by-step") helps Mistral but slightly harms ChatGPT and ChatGPT-Instruct with DIO-top8 Hwang et al. (2023). On the other hand, SC improves all models. Therefore, we attribute CoT’s limitation to the inconsistency of its explanations (3). Meanwhile, ChOiRe with CoO consistently attains improvements, verifying the effectiveness of explicitly requiring the model to analyze all the personae. Finally, ChatGPT, ChatGPT-Instruct, and Mistral show improvements by selecting only $4.79/12$ and $5.59/12$, $8.83/12$ explicit personae on average, respectively. This suggests that over half of explicit personae may be noisy for models to predict opinions.
Fine-grained Prompting Results.
Diving deeper into the benchmark topics in table 5, ChOiRe achieves SOTA results in $8/15$, $8/15$, $11/15$, $13/15$ topics for ChatGPT and ChatGPT-Instruct, GPT-4, and Mistral. The improvements are especially huge for some topics. For example, compared with the best among baselines, it improves GPT-4 up to $12.08\%$ accuracy on Views on gender, ChatGPT up to $9.82\%$ on Economic Inequality. We also specifically compare ChOiRe with the best baseline DIO-top8 + SC in section B.3, showing $8/12$ improvements for ChatGPT and ChatGPT-Instruct. We further plot the accuracy distribution over users of ChOiRe, specifically for ChatGPT in fig. 4. We see that the majority accuracy is $0.5$, with a few users scoring zero and over $20$ achieving perfection.
Fine-tuning Results.
Table 2 presents our fine-tuning outcomes. Notably, leveraging the ChOiRe’s FEA and LLMtop-$K$ steps on the fine-tuning data yields substantial enhancements for GPT-2-large and FlanT5-large, showcasing relative accuracy improvements of $18.44\%$ and $7.38\%$ respectively. Remarkably, ChOiRe’s FEA and LLMtop-$K$ steps bring FlanT5-large’s performance on par with GPT-4, despite GPT-4’s significantly stronger capability. Furthermore, ChOiRe’s LLMtop-$K$ proves particularly beneficial for enhancing FlanT5-base. Surprisingly, GPT-2-base performs well even without user demographic and ideological information, possibly due to potential contamination Sainz et al. (2023) with public polling data from OpinionQA.
Human Evaluation Results.
Model FEA Satis. LLMtopK Satis. Rea. Foll. Inst.
[…]
Our human evaluation results in table 3 reveal three key findings. First, ChatGPT and ChatGPT-Instruct achieve similar performance in filtering explicit personae and ranking opinions, while Mistral achieves lower results. While ChatGPT excels slightly in ranking, ChatGPT-Instruct performs slightly better in explicit personae selection. Three models proficiently filter unnecessary explicit personae, but ranking opinions poses a more challenging task intuitively and empirically, with a common error being the inconsistent relevance ranking of opinions, sometimes misplacing high-level relevance. Second, four models effectively generate reasonable thoughts leading to the final answer, and GPT-4 performs the best. Finally, ChatGPT and ChatGPT-Instruct follow our instructions to explain and analyze the explicit and implicit personae provided one by one with CoO significantly better than GPT-4 and Mistral, achieving nearly perfect scores of $3$. We hypothesize that this is because ChatGPT and ChatGPT-Instruct excel in following instructions, while GPT-4 is optimized for completing texts.
6 Discussion
We discuss the main analyses in this section. Extra analyses are presented in appendix B.
6.1 Methodology Analysis
Ablation of FEA.
To gauge the impact of filtering unnecessary explicit personae (FEA) on performance, we experiment with applying FEA exclusively to the baseline DIO-top8 Hwang et al. (2023), denoted as DIO-top8 + FEA in table 1. The results indicate enhancements with DIO-top8 + FEA achieving a 0.8%, 1.3%, 1.9% accuracy performance boost on ChatGPT, ChatGPT-Instruct, and Mistral respectively. This underscores the effectiveness of eliminating irrelevant explicit personae in improving the models’ ability to understand and predict human opinions.
FEA via Topics.
To understand the explicit personae filtered by LLMs across various topics, we document the top-3 removed personae in section B.5. We observe that "Citizenship" is consistently the most frequently removed attribute, followed by "Race". This could be due to LLMs treating these as sensitive information, prioritizing respect and unbiased text generation. Another explanation may be the lack of correlation between citizenship/race and opinions in the US-centric OpinionQA dataset. Additionally, we also see that ChatGPT often categorizes ‘‘Marital status" as non-useful, ChatGPT-Instruct commonly removes ‘‘Frequency of religious attendance", and ‘‘Gender" got removed by Mistral, revealing potential biases in LLMs.
LLMtop-$K$ versus Top-$K$.
From table 1, DIO-LLMtop8 outperforms DIO-top8 by $1.6\%$, $1.6\%$, $3.8\%$ accuracy on ChatGPT, ChatGPT-Instruct, Mistral confirming that prioritizing meaning and usefulness improves opinion prediction. One possible explanation for this can be the orders ranked by semantic similarity scores only consider ranking with the input questions Hwang et al. (2023), while our orders consider both input questions and their answer choices (fig. 1). We further explore two key aspects: (1) The agreement of LLM-orders and semantic similarity orders, and (2) Points of maximum disagreement between these orders. To measure the ranking agreements, we calculate Kendall’s Tau correlation coefficient Kendall (1938) between the orders generated by ChatGPT, ChatGPT-Instruct, and Mistrial and orders sorted by semantic similarity scores, and the results are presented in fig. 6 and fig. 7. Surprisingly, for ChatGPT and ChatGPT-Instruct, we find that the two ranking orders have minimal monotonous relations with means approximating 00 and low standard deviations showing no agreement. For Mistral, we find a low agreement with a mean of $0.43$ score. These low and no agreements further verify that ranking by usefulness can be very different from ranking by semantic similarity. We also deep dive into cases with notable order variations to address (2). Section E.2 illustrates one such case in the "Guns" topic. We derive three observations. First, not all top-8 opinions by semantic similarity scores help predict the opinion. For example, the $16$-th opinion, despite having a relatively high semantic similarity score with the question which might offer some perspective on the prevalence of guns in the user’s community during the upbringing, is less directly related to the question. This is similar to the $18$-th opinion which is also less relevant. Meanwhile, several important opinions are deselected by the semantic-similarity-based method, such as the $6,3,4,10$-th ones, which are chosen by the LLM. The $6$-th one is critical, and directly relevant because it assesses the person’s attitude toward safety measures related to gun ownership. Finally, by using LLMtop-$K$ order, the model predicts the opinion accurately, whereas the semantic similarity order leads to an incorrect prediction.
Opinions Order Analysis of LLMtop-$K$ Step.
The performance difference of DIO-top8 and DIO-LLMtop8 in table 1 highlights that LLMs are sensitive to the chosen implicit personae opinions. An important question arises: Are LLMs also affected by the input order of implicit persona opinions in the ranking step (section 3)? Our discovery confirms sensitivity, but with reasonable overlap when $K$ is sufficiently large. We randomly select $300$ questions, shuffle implicit persona opinions four times with different seeds, and record four LLM ranking outputs for each. We also collect one more LLM ranking output by feeding implicit personae opinions in semantic similarity order. For each $K\in\{1,2,...,20\}$, we calculate the pairwise Overlap coefficient Vijaymeena and Kavitha (2016) among the five ranking outputs, averaging them as the LLM ranking consistency score for each $K$. The scores, shown in fig. 5, indicate that for $K\geq 8$, the ranking outputs overlap well with a score of .$\geq.6$ for both models. Despite this, is there substantial variance in model performance across random seeds? Our findings reveal no significant variance, with the variants statistically outperforming the baseline DIO-top8. Specifically, we assess ChatGPT and Mistral with DIO-LLMtop8 on $3$ out of $4$ random seeds, detailed in section B.6. The results demonstrate relatively small standard deviations in their performance, and critical values of $99\%$ CI of DIO-LLMtop8 under t-test for both models surpass DIO-top8, confirming that LLMtop8’s effectiveness is not due to randomness.
CoO versus CoT.
Figure 2: Consistency scores of the baseline DIO-top8 (ChatGPT) with CoO and CoT.
Table 1 indicates that Chain-of-Thought (CoT) Kojima et al. (2022) slightly harms baseline DIO-top8 performance for ChatGPT and ChatGPT-Instruct. Conversely, our Chain-of-Opinion reasoning (CoO) enhances overall performance for all models. To investigate the consistency of CoT and CoO, we design an experiment with ChatGPT, DIO-top8 where we randomly select $100$ question-answer pairs and sample $5$ answers per pair using CoT and CoO, at $3$ different temperatures $0.3$, $0.6$, $0.9$. For each prompting technique, we measure the percentage of questions that all $5$ answers sampled have the same result, as the consistency score. The results are illustrated in fig. 2 showing that CoO brings slightly better consistent answers compared to CoT, especially when the temperature is high which verifies CoO potentially enhances the reliability of LLMs.
Dynamic Numbers of Opinions Analysis.
Table 4 illustrates our analysis answering two research questions: (1) How frequent can’t LLMs answer the question? and (2) How do LLMs perform when more opinions than $K=8$ are provided in ChOiRe? Our findings show that, firstly, with $8$ opinions, GPT-4 exhibits the highest percentage of unanswered questions, while Mistral answers all the questions. Secondly, increasing the number of opinions beyond $8$ reduces this percentage across models, confirming our hypothesis regarding the lack of implicit personae opinions when fixing $K=8$ in section 3. Lastly, while including more opinions could possibly harm the performance of models, our answer consistency strategy enables LLMs to achieve the best results across three different K values.
Model ChatGPT ChatGPT-Inst GPT-4 Mistral
% of ITA of DIO-LLMtop8 + FEA + CoO 0.61 1.32 9.71 0.00
DIO-LLMtop8 + FEA + CoO 51.90 53.01 59.02 54.21
% of ITA of DIO-LLMtop10 + FEA + CoO 0.12 1.01 5.44 0.00
DIO-LLMtop10 + FEA + CoO 51.55 52.74 58.88 53.88
% of ITA of DIO-LLMtop12 + FEA + CoO 0.00 0.66 3.12 0.00
DIO-LLMtop12 + FEA + CoO 51.60 52.31 59.11 52.96
ChOiRe 52.21 53.26 59.30 54.43
Table 4: Extra analysis on ChatGPT, ChatGPT-Instruct, GPT-4, and Mistral. ITA stands for "Impossible To Answer".
6.2 Error Analysis
FEA Misses Key Explicit Personae.
Despite showing promising results in removing unuseful explicit personae depicted in table 3, we observe that LLMs sometimes misselect relevant personae. One such example is the top-left of section E.6. We observe that in this case, our annotators can’t grade a high FEA satisfaction score because "Education" and "Age" are also two important personae as they can influence one’s understanding of workplace dynamics significantly, which are deselected by ChatGPT.
LLMtop-$K$ Opinions Include Less Relevant Ones.
While LLMs generally demonstrate a commendable ability to rank implicit opinions by usefulness, as exemplified in section E.2, we also observe they frequently include less relevant, or even irrelevant opinions to the ranked list such as in section E.6-bottom. We attribute this to the challenge of this task, even for humans it might require substantial cognitive effort.
LLMs May Not Follow the Instructions.
Although ChatGPT and ChatGPT-Instruct demonstrate a robust ability to adhere to our instructions for opinion prediction via CoO, the same level of proficiency is not observed in Mistral and GPT-4, as shown in section E.6-top-right. We posit this disparity arises from the fact that ChatGPT and ChatGPT-Instruct excel in comprehending and executing human instructions, while GPT-4 excels primarily in generating coherent text.
7 Conclusion
We propose ChOiRe, a four-step solution framework for individual opinion prediction via differentiating the utilization of user’s explicit versus implicit personae. We further introduce Chain-of-opinion reasoning and answer consistency over variable numbers of input implicit personae guiding the models to derive thorough predictions. ChOiRe achieves strong SOTA results with limited inference calls, demonstrating its strong effectiveness. Additionally, Steps (i) and (ii) of ChOiRe significantly improve the fine-tuning of opinion-aligned models. We strongly suggest that our method should only be used for positive moral intents, avoiding making LLMs echo chambers Vicario et al. (2016). In the future, we will focus on developing frameworks that utilize explicit and implicit personae more efficiently.
Limitations
One limitation of our proposed ChOiRe framework is that it requires the LLMs to have a good capability in following human instructions to solve tasks such as selecting explicit personae, ranking historical opinions, and explaining personae and opinions one by one via CoO. However, we foresee that this limitation is going to be overcome by cutting-edge AI language models, in the present and near future. Additionally, our method also utilizes user’s personal information from explicit and implicit personae, which may be sensitive to some audiences and not be fully available in the real world. However, to what extent is the personal information provided, our ChOiRe is still able to offer reasonable opinion predictions since it is not constrained by the number of provided explicit personae, or the number of user historical opinions.
Ethical Considerations
Characterizing and predicting human opinions with LLMs can be directly applied to personalize and align machines to users’ values, and cultural beliefs. Nonetheless, there exist unwanted situations when LLMs with our techniques can be misused for unethical purposes and biased opinions.
Bias Amplification and Fairness.
A personalized LLM allows users to reinforce their existing beliefs and potentially amplify biased or unethical perspectives, leading to the creation of echo chambers Vicario et al. (2016). This can ultimately harm users by reinforcing polarized or undesirable views. To mitigate this issue, the Chain-of-Opinion (CoO) reasoning from our proposed ChOiRe involves presenting user demography or ideology group responses alongside personalized answers. Additionally, CoO can encourage users to reflect on their previous viewpoints.
Privacy and Consent.
Users may not always be aware of or have control over the extent of personalization applied to the content they receive. Therefore, empowering users to have control over AI-generated opinions is essential. Users should be able to customize and adjust the explicit and implicit personae used for opinion prediction. This customization can help mitigate potential biases and provide individuals with AI-generated opinions that align more closely with their values and preferences.
Human Evaluation.
Through human evaluations, we observe that our proposed method does not generate any discriminatory, insulting responses. We validate the intermediate steps of our proposed ChOiRe by human evaluation which involves manual labor. We hire annotators to score, and the hourly pay is set to \$$\$15$, which is higher than the local statutory minimum wage. Therefore, we do not anticipate any major ethical concerns raising from human evaluations.
Appendix A Baselines
A.1 Baselines Implementation Details
Prompting.
[…]
Model ChatGPT/ChatGPT-Inst/GPT-4/Mistral
Guns Auto. & driverless vehicles Views on gender Com. types & sex. harassment Race
W/o persona 53.07/37.30/——-/30.48 47.73/48.26/——-/41.72 50.53/42.94/——-/37.39 47.73/41.67/——-/29.34 41.95/45.28/——-/37.55
DIO-top8 53.87/57.00/60.39/44.73 45.33/44.78/53.22/41.72 53.21/52.15/63.73/40.09 43.47/45.24/42.86/35.45 43.06/44.65/55.17/41.11
DIO-top8 + CoT 54.55/52.33/——-/55.48 47.22/46.77/——-/49.00 48.11/57.67/——-/54.28 42.39/42.26/——-/42.01 45.63/43.40/——-/49.78
DIO-top8 + SC 54.40/52.85/——-/56.57 43.73/48.26/——-/52.31 55.61/56.44/——-/56.30 45.33/40.48/——-/42.01 45.00/43.40/——-/50.00
ChOiRe 57.06/58.21/63.37/58.00 49.25/51.92/50.00/53.75 59.23/53.07/71.43/57.78 39.88/44.14/47.96/42.08 42.77/47.28/50.57/51.44
Gender & Leadership America in 2050 Trust in science Biomedical & food issues Misinformation
W/o persona 53.13/50.83/——-/43.51 39.73/39.13/——-/41.95 50.40/47.29/——-/48.34 53.87/53.63/——-/53.21 46.93/40.38/——-/53.63
DIO-top8 48.27/54.70/65.55/50.23 46.93/46.20/43.70/35.14 54.93/61.58/61.54/51.65 52.27/55.86/58.03/52.78 49.33/52.11/52.71/50.77
DIO-top8 + CoT 48.58/50.83/——-/55.79 43.05/48.91/——-/43.76 54.10/65.02/——-/58.28 56.91/57.54/——-/57.08 49.57/53.99/——-/53.19
DIO-top8 + SC 49.07/53.60/——-/57.87 45.87/47.83/——-/46.03 56.27/65.52/——-/58.94 53.07/57.54/——-/58.58 45.00/53.52/——-/53.85
ChOiRe 52.22/57.78/63.03/57.87 49.46/48.99/45.37/47.50 56.43/55.50/68.46/60.37 54.75/57.26/61.61/58.58 46.45/53.62/57.36/53.85
Privacy & Surveilance Family & Relationships Economic inequality Global attitudes Political views
W/o persona 43.24/40.28/——-/33.64 47.06/44.36/——-/46.08 43.67/49.15/——-/34.07 46.13/46.71/——-/40.42 40.80/48.95/——-/46.20
DIO-top8 53.24/47.22/47.73/43.31 57.22/57.89/62.50/47.42 45.60/51.98/63.81/41.87 49.60/57.23/66.67/41.27 56.80/46.85/62.07/44.13
DIO-top8 + CoT 53.38/47.22/——-/56.91 59.57/55.64/——-/54.36 47.65/51.98/——-/51.45 46.42/56.58/——-/51.06 53.30/45.45/——-/50.80
DIO-top8 + SC 54.05/47.22/——-/58.06 55.35/54.89/——-/57.04 46.13/51.98/——-/52.76 46.42/55.26/——-/52.89 57.33/47.55/——-/51.67
ChOiRe 54.29/53.33/52.27/58.06 60.00/58.77/63.89/58.50 52.33/50.13/64.76/51.89 44.74/55.26/64.58/52.76 51.05/53.74/67.82/53.34
Table 5: Fine-grained accuracy results of ChatGPT/ChatGPT-Instruct/GPT-4/Mistral. DIO stands for Demographic + Ideology + Opinions (section 4).
Fine-tuning.
[…]
Appendix B Extra Analysis
B.1 Additional Baseline Comparisons
Model ChatGPT Mistral-7B-Instruct-v0.2
DIO-top8 50.22 44.16
DIO-top8 + FEA 50.64 44.99
DIO-top8 + Random FEA (S=2000) 49.47 42.23
DIO-top8 + Random FEA (S=2024) 48.85 43.36
DIO-LLMtop8 51.03 45.86
DIO + Random LLMtop8 (S=2000) 48.13 44.58
DIO + Random LLMtop8 (S=2024) 49.21 43.84
Table 6: Accuracy results of ChatGPT and Mistral with two trivial variants with two different random seeds 2000 and 2024 in section B.1.
In this section, we compare ChOiRe’s FEA and LLMtop-$K$ steps with two simple variants outlined in table 6. Given ChatGPT and Mistral’s strong performance with just $4.79/12$ and $8.83/12$ explicit persona attributes, a crucial question arises: Can comparable performance be achieved by randomly selecting $5/12$ and $9/12$ explicit persona attributes instead of relying on LLMs? The first variant, DIO-top8 + Random FEA, involves randomly selecting $5/12$ and $9/12$ explicit persona attributes. The second variant entails randomly selecting $8$ implicit persona opinions instead of using ChOiRe’s LLMtop-$K$ step. From table 6, we find that randomly selecting explicit persona attributes significantly harms the performance of both models due to the removal of important attributes. Additionally, randomly selecting $8$ implicit persona opinions also adversely affects the models, particularly ChatGPT. These observations underscore the effectiveness and importance of ChOiRe’s FEA and LLMtop-$K$ steps.
B.2 Fine-grained Results of ChatGPT, ChatGPT-Instruct, and GPT-4
Table 5 presents the fine-grained results of ChOiRe and baselines for ChatGPT, ChatGPT-Instruct, and GPT-4.
B.3 ChOiRe versus Self-Consistency
Fig. 3 presents the improvements per topic comparision between ChOiRe and the SOTA baseline with Self-Consistency Wang et al. (2023b) DIO-top8 + SC. We observe that ChOiRe improves $11/15$ topics for both ChatGPT and ChatGPT-Instruct.
Figure 3: % of improvements over the SOTA method (DIO-top8 + SC) with ChatGPT-Instruct (left) and ChatGPT (right).
B.4 Accuracy Distribution over Users
Fig. 4 shows the accuracy distribution over users of ChOiRe with ChatGPT. We see that the peak accuracy is at 0.5 for the majority, with a few users scoring zero and over 20 achieving perfection.
Figure 4: Frequency distribution of accuracy over users by ChOiRe.
B.5 Top-3 Removed Explicit Personae Attributes
Table 7 presents the top-3 explicit personae that got removed the most by the LLMs. Among the removed personae, "Citizenship" appears to be the highest-frequency one across models, followed by "Race".
Topic ChatGPT ChatGPT-Instruct Mistral-7B-Instruct-v0.2
Guns ’Citizenship’, ’Race’, ’Marital status’ ’Citizenship’, ’Frequency of religious attendance’, ’Religion’ ’Citizenship’, ’Education’, ’Religion’
Automation & driverless vehicles ’Citizenship’, ’Race’, ’Marital status’ ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Citizenship’, ’Religion’, ’Frequency of religious attendance’
Views on gender ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Citizenship’, ’Religion’, ’Frequency of religious attendance’
Community types & sexual harassment ’Citizenship’, ’Race’, ’Gender’ ’Citizenship’, ’Frequency of religious attendance’, ’Race’ ’Education’, ’Race’, ’Political Party’
Biomedical & food issues ’Citizenship’, ’Race’, ’Marital status ’Citizenship’, ’Race’, ’Marital status’ ’Citizenship’, ’Race’, ’Marital status’
Gender & Leadership ’Citizenship’, ’Race’, ’Region’ ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Region’, ’Race’, ’Citizenship’
America in 2050 ’Citizenship’, ’Race’, ’Marital status’ ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Citizenship’, ’Frequency of religious attendance’, ’Race’
Trust in science ’Citizenship’, ’Marital status’, ’Race’ ’Citizenship’, ’Race’, ’Marital status’ ’Citizenship’, ’Race’, ’Region’
Race ’Citizenship’, ’Marital status’, ’Age’ ’Citizenship’, ’Age’, ’Religion’ ’Marital status’, ’Education’, ’Age’
Misinformation ’Citizenship’, ’Marital status’, ’Race’ ’Citizenship’, ’Marital status’, ’Race’ ’Citizenship’, ’Race’, ’Religion’
Privacy & Surveillance ’Citizenship’, ’Race’, ’Marital status’ ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Religion’, ’Race’, ’Region’
Family & Relationships ’Citizenship’, ’Race’, ’Region’ ’Citizenship’, ’Race’, ’Frequency of religious attendance’ ’Citizenship’, ’Race’, ’Religion’
Economic inequality ’Citizenship’, ’Frequency of religious attendance’, ’Race’ ’Citizenship’, ’Frequency of religious attendance’, ’Race’ ’Gender’, ’Citizenship’, ’Religion’
Global attitudes ’Marital status’, ’Race’, ’Citizenship’ ’Citizenship’, ’Marital status’, ’Race’ ’Gender’, ’Frequency of religious attendance’, ’Marital status’
Political views ’Citizenship’, ’Marital status’, ’Frequency of religious attendance’ ’Citizenship’, ’Frequency of religious attendance’, ’Race’ ’Frequency of religious attendance’, ’Gender’, ’Citizenship’
Table 7: Top-3 explicit personae that got removed the most by the LLMs.
B.6 Ranking Consistency for LLMtop-$K$ Step
We record the average Overlap coefficient Vijaymeena and Kavitha (2016) among $5$ ranking outputs from $5$ input strategies in fig. 5. The performance of those input strategies is further presented in table 8 on $300$ random samples.
Figure 5: ChatGPT and Mistral-7B-Instruct-v.02 overlap coefficient values for different values of $K$. We observe that for $K$ is large enough ($K\geq 8$), the coefficient value is relatively acceptable ($\geq 0.6$).
Model Method Semantic Similarity Order Seed = 2024 Seed = 5 Seed = 2000 Seed = 15 Std
ChatGPT DIO-LLMtop8 - 51.03 50.95 51.11 - 0.0652
Mistral-7B-Instruct-v0.2 DIO-LLMtop8 - 45.86 45.55 45.36 - 0.2060
Table 8: Accuracy results of ChatGPT and Mistral on our test set with DIO-LLMtop8 where different orders of input implicit persona opinions are tested for LLMtop-K step.
B.7 Kendall’s Tau Scores for Ranking Agreements
Figure 6: Left: Ranking agreements between ChatGPT top-$K$ and semantic similarity top-$K$. Right: Between ChatGPT-Instruct top-$K$ and semantic similarity top-$K$. One example that has a high disagreement score is shown in section E.2.
Figure 7: Ranking agreements between Mistral top-$K$ and semantic similarity top-$K$.
Fig. 6 shows our ranking agreement scores between ChatGPT and Semantic similarity metric (Left), and ChatGPT-Instruct and Semantic similarity metric (Right). We observe that the two ranking orders have minimal monotonous relations with means approximating 0 and low standard deviations. More specifically, with ChatGPT, the maximum agreement is 0.6000 while the minimum is -0.5895 and the Kurtosis is -0.2173. For ChatGPT-Instruct, the maximum is slightly lower with 0.5473, while the minimum is -0.7368 which is smaller ChatGPT, and the Kurtosis is -0.1017.
B.8 Consistency Scores
Table 9 presents the exact consistency scores for the fig. 2. Besides CoO consistently outperforming CoT, we also observe that when the temperature is increased, the consistency score is decreased, which is intuitive.
Model Temperature Consistency Score (%)
DIO-top8 + CoT 0.3 84
DIO-top8 + CoO 0.3 86
DIO-top8 + CoT 0.6 79
DIO-top8 + CoO 0.6 82
DIO-top8 + CoT 0.9 58
DIO-top8 + CoO 0.9 60
Table 9: Consistency scores of CoT and CoO on $100$ random question-answer pairs. We sample $5$ answers per question and measure the % of questions that have all $5$ identical answers.
Appendix C Prompts and Prompts Analysis
C.1 Prompt Templates for Filtering Explicit Personae
We present the prompt template for selecting relevant explicit personae for answering the question below. The template is hand-crafted and we use Chain-of-Thought (CoT) prompting Kojima et al. (2022) via adding "answer the above question step by step".
A person can be described by the following attributes: {original_attribute_list} Based on the above list of demographic information above, now I give you a new question with possible answer choices: Question: ’{test_question}’ Answer choices: ’{test_choices}’ Please analyze which attributes in the demographic information are useful for you to answer the above question step by step. Give me the output in the Python list format: [...] Give me the answer in the format below: Explanations: ... Answer: [...]
C.2 Prompt Templates for Implicit Feature Ranking
We provide our hand-crafted prompt template for ranking implicit personae opinions in the usefulness order below:
Given social behavior question-answer pairs answered by a user about his/her opinions about {subtopic}: {original_persona_question_order} You are an expert in analyzing the social behaviors of a user. Given a new question asking him/her: ’{test_question}’ Your task is to sort the list of given question-answer pairs in descending order such that the first question-answer pair brings the most useful information to answer the new question, whilst the last question-answer pair brings the least useful information. Give me the answer in the form of a Python list of indexes: Answer: [...]
C.3 Prompt Templates for Baselines Techniques
We use the same prompt templates for ChatGPT OpenAI (2022), ChatGPT-Instruct OpenAI (2023a), GPT-4 OpenAI (2023b). The template prompts for baselines are presented below.
C.3.1 W/o Persona Santurkar et al. (2023)
The W/o Persona prompt is provided below.
Question: {question} Answer choices: {choice} Complete the answer by the following format without any explanation: Answer: A. or B. or C. or D. or E...
C.3.2 DIO-top8 Hwang et al. (2023)
The DIO-top8 prompt is provided below.
A person can be described as follows: {explicit_persona_str} The person has the following opinions on {topic}. Opinions: {implicit_persona_str} Based on the above information, which answer choice is the user most likely to choose? Question: {question} Answer choices: {choice} Give the answer in the format: Answer: A. or B. or C. or D. or E....
C.3.3 Self-refine Madaan et al. (2023)
The Self-refine prompts Madaan et al. (2023) are provided below, feedback step and refine step respectively.
You are given a question and an answer for that question. Analyze the question and the answer and provide some feedback on the answer to the question. Don’t change the answer, just provide feedback. Question: {test_question} Choices: {choices} Answer: {selected_choice} Feedback:
You are given a question, an answer to that question and a feedback to the answer. Based on the feedback, refine your answer and generate the final answer in around 170 words. Question: {test_question} Answer: {selected_choice} Feedback: {feedback} Refined answer: new_choice + explanation
C.3.4 Chain-of-Thought Kojima et al. (2022)
The CoT prompt template is provided below.
A person can be described as follows: {explicit_persona_str} The person has the following opinions on {topic}. Opinions: {implicit_persona_str} Based on the above information, answer the following question step-by-step: Question: {question} Answer choices: {choice} Give the answer in the format: Answer: A. or B. or C. or D. or E.... Explanations:...
C.3.5 Chain-of-Opinion (Ours)
Our CoO prompt template is provided below.
A person can be described as follows: {explicit_persona_str} The person has the following opinions on {topic}. Opinions: {implicit_persona_str} Based on the above information, answer the following question step-by-step by explaining and analyzing each of the Opinions and Demographic Information: Question: {question} Answer choices: {choice} Give the answer in the format: Answer: A. or B. or C. or D. or E.... Explanations:...
C.4 Prompting Costs for API Models
Our prompting costs for API models are reported in table 10. We observe that for ChatGPT and ChatGPT-Instruct, ChOiRe costs around $7$ and $10$ more US\$ dollars in total compared to baseline DIO-top8 + SC. However, these extra amounts of costs are worth it because we gain significant improvements over all the baselines and especially huge improvements for some topics. Additionally, for GPT-4, ChOiRe costs a similar price with the baseline DIO-top8 while DIO-top8 + SC costs nearly double our price. This is because we perform the FEA and LLMtop-$K$ steps of ChOiRe by ChatGPT, which are relatively cheap.
DIO-top8 DIO-top8 + CoT DIO-top8 + SC ChOiRe Model
Ave. consumed #tokens 562.72 623.62 995.89 3142.86 ChatGPT
Total US\$ 3.01 3.73 6.82 13.95 ChatGPT
Ave. consumed #tokens 562.72 630.58 1019.31 3121.72 ChatGPT-Instruct
Total US\$ 3.12 3.84 7.11 19.99 ChatGPT-Instruct
Ave. consumed #tokens 559.27 - 1021.14* 3180.82 GPT-4
Total US\$ 91.19 - 226.15* 123.30 GPT-4
Table 10: Prompting cost analysis of ChOiRe and other baselines as of 1st Feb 2024. * denotes our estimation on $50$ random samples.
Appendix D Human Evaluation
D.1 Human Rating Instructions
Our details of human rating instructions are provided in table 11 for all the criteria. It is worth noting that selecting all features can’t get a high FEA Satisfaction score, according to our instructions. In addition, if the selected explicit personae fall among several scores, the annotators are instructed to take the minimum score.
Criterion Scoring Instruction
1: The number of filtered-out explicit personae that are directly relevant for answering the question is more than 3.
1: The number of selected explicit personae that are somewhat irrelevant for answering the question is more than 3.
2: The number of filtered-out explicit personae that are directly relevant for answering the question is 2 or 3.
FEA Satisfaction 2: The number of selected explicit personae that are somewhat irrelevant for answering the question is 2 or 3.
3: The number of filtered-out explicit personae that are directly relevant for answering the question is less than or equal to 1.
3: The number of selected explicit personae that are somewhat irrelevant for answering the question is less than 2.
1: Among the top-8 implicit persona opinions, the number of less relevant opinions for answering the question is more than 4.
LLMtop-$K$ Satisfaction 2: Among the top-8 implicit persona opinions, the number of less relevant opinions for answering the question from 2 to 4.
3: Among the top-8 implicit persona opinions, the number of less relevant opinions for answering the question is less than or equal to 1.
1: The CoO has limited or flawed reasoning thoughts with inadequate support.
CoO Reasonableness 2: The CoO has some reasoning thoughts with decent support but room for improvement.
3: The CoO has strong, clear, and well-supported reasoning thoughts with a comprehensive understanding.
1: The generated CoO explanation does not mention more than $4$ attributes/opinions from explicit and implicit personae.
CoO Follow the Instruction 2: The generated CoO explanation somewhat follows the instruction by involving more than $4$ attributes/opinions but room for improvement.
3: The generated CoO explanation follows perfectly the instruction via explaining all the explicit and implicit attributes one by one.
Table 11: Human rating instructions. FEA, LLMtop-$K$, and CoO stand for Filtering Explicit Personae Attributes, Implicit Personae Opinions Ranking, and Chain-of-Opinion reasoning (section 3).
Appendix E Examples
E.1 FEA Example with ChatGPT
Fig. 8 shows an FEA example with ChatGPT. We observe that by removing unnecessary explicit personae including "Age", "Citizenship", "Education", "Income", "Marital Status", "Race", "Frequency of religious attendance", ChatGPT predicts the opinion accurately, while without removing, an incorrect prediction was made.
Figure 8: FEA example with ChatGPT.
E.2 Example of High Disagreement between Rankings
Fig. 9 illustrates one example of the high disagreement between orders by semantic similarity scores and LLM (ChatGPT). We derive three observations, as discussed in section 6.1. First, not all top-8 opinions by semantic similarity scores help predict the opinion. For example, $16$-th opinion, despite having a relatively high semantic similarity score with the question which might offer some perspective on the prevalence of guns in the user’s community during the upbringing, is less directly related to the question. This is similar to the $18$-th opinion which is also less relevant. Meanwhile, several important opinions are deselected by the semantic-similarity-based method, such as the $6,3,4,10$-th ones, which are chosen by the LLM. The $6$-th one is critical, and directly relevant because it assesses the person’s attitude toward safety measures related to gun ownership. Finally, by using LLMtop-$K$ order, the model predicts the opinion accurately, while an incorrect prediction is made with the semantic similarity order.
Figure 9: Example of the high disagreement between orders by semantic similarity scores and LLM (ChatGPT).
E.3 Example of Inconsistent Answers Generated by CoT
Fig. 10 illustrates an example of the inconsistent answers generated by ChatGPT with Chain-of-Thought Kojima et al. (2022) (CoT). It is observed that different subsets of top-8 implicit personae opinions are mentioned in the two explanations, leading to varied final answers.
Figure 10: Example of the inconsistent answers generated by ChatGPT with Chain-of-Thought.
E.4 Example of Chain of Opinion Reasoning
Fig. 11 presents an example of the answer generated by ChatGPT using Chain of Opinion (ours) versus Chain of Thought Wei et al. (2022) prompting methods.
Figure 11: Example of an answer generated by Chain of Opinion versus Chain of Thought prompting with ChatGPT.
E.5 Example of Answer Consistency with Dynamic Numbers of Opinions
Fig. 12 shows an example of the answer generated by GPT-4 using Chain of Opinion (ours) reasoning with different numbers of provided historical opinions.
Figure 12: Example of our answer consistency technique (ours), generated by GPT-4.
E.6 Error Analysis Examples
Fig. 13 illustrates our error analysis examples of ChOiRe with ChatGPT. The top-left frame is an example of FEA missing key explicit personae. The bottom one is an instance demonstrating the error of the LLMtop-$K$ algorithm including less relevant opinions. The top-right rectangular is an example from GPT-4, showing that it does not follow human instructions to predict opinion via chain-of-opinion reasoning.
Figure 13: Error analysis examples of ChOiRe with ChatGPT.